OpenStudy (anonymous):

Find the formula for another line passing through(1,7) and (2,-2)

OpenStudy (anonymous):

Use point slope formula (y2-y1) / (x2-x1)

OpenStudy (anonymous):

so its (-2 - 7) / (2-1) = -9 / 1 = -9 is your answer for slope

OpenStudy (anonymous):

The first part is not point slope formula

OpenStudy (anonymous):

It's the formula to get the slope between two points

OpenStudy (anonymous):

point slope is y-y1=m(x-x1)

OpenStudy (anonymous):

where (x1,y1) is one of your points while m is your slope. @ilikephysics2 showed you how to get the slope and you are given two points in the question so use either one and you are set. PLug and enjoy your answer.